OpenStudy (anonymous):

Could somebody check whether i solved the question correctly? If the straight lines 3x-y=-2 and ax+2y=3 are parallel then a=?

OpenStudy (anonymous):

Y=3x+2 Y=-ax/2 +3/2

OpenStudy (anonymous):

The both must have a common slope since they are parallel

OpenStudy (anonymous):

They*

OpenStudy (anonymous):

3x = -ax/2

OpenStudy (anonymous):

I get a= -6

OpenStudy (nikato):

u can just get rid of the x and solve for a, so 3=-a/2

OpenStudy (nikato):

yes, a=-6 is correct

OpenStudy (anonymous):

Is there another way of solving it?

OpenStudy (anonymous):

By using the elimination method?

OpenStudy (nikato):

i dont think so.

OpenStudy (nikato):

elimination would help u find the x and y, but in ur case u were trying to find a coefficient,a